# Energy, enstrophy and helicity transfers in polymeric turbulence

### v1

@@ 5.1 Scale-by-scale budget
$\omega_{r}\sim\frac{u_{r}}{r}\sim\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle^{1/3}r^{-2/3},$ (24)
@@
which becomes maximum as $r\to\eta$, so that $\omega_{\eta}\sim\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle^{1/3}\eta^{-2/3}$. Now, to estimate the cumulative vortex stretching $\left\langle{V_{st}}\right\rangle$ up to scale $r$,
@@
$\left\langle{V_{st}}\right\rangle(r)$ $=\frac{1}{S(r)}\int_{V(r)}\text{d}V\ 2\left(\left\langle{\delta\omega_{i}\omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)}\right\rangle+\left\langle{\delta\omega_{i}\delta\omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}}\right\rangle\right),$ (25)
we begin by estimating the scale-by-scale vortex stretching $v^{st}_{r}$. This is given by the terms within the integral sign in the above equation 25, i.e.,
$v^{st}_{r}\sim\omega_{r}\omega_{r}\frac{u_{\eta}}{\eta}.$ (26)
@@
Indeed, $v^{st}_{r}$ also gets a maximal contribution from the smallest scales (see also Davidson et al., 2008) which is given as
$v^{st}_{\eta}\sim\omega^{2}_{\eta}\ \frac{u_{\eta}}{\eta}\sim\left(\frac{\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle^{1/3}}{\eta^{2/3}}\right)^{3}\sim\frac{\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle}{\eta^{2}}\sim\left\langle{\varepsilon_{f}^{\delta\omega^{2}}}\right\rangle,$ (27)
@@
where we have used the $\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle\sim\langle\varepsilon_{f}^{\delta q^{2}}\rangle/\eta^{2}$ dimensional estimate. Ignoring the contribution to vortex stretching from larger scales, we can estimate the cumulative vortex stretching upto a scale $r$ as $\langle V_{st}\rangle(r)\approx 1/S(r)\int_{V(r)}\text{d}Vv^{st}_{\eta}\sim r^{-2}r^{3}\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle\sim\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle r$, which is consistent with the data plotted in figure 9. Moreover, the nonlinear flux shows a $\langle\Phi_{f}^{\delta\omega^{2}}\rangle(r)\sim r^{-1}$ scaling and its magnitude increases as $r$ decreases within the inertial range. Since enstrophy is mainly generated at the small scales, its transfer rate among scales strengthens as $r$ decreases, with a maximum close to the transition from the inertial to the dissipative range of scales. We follow Davidson et al. (2008) to use dimensional analysis to explain the observed scaling law for $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$. The idea is that in the inertial range of scales the enstrophy cascade is based on mechanisms that are similar to vortex stretching. For example, let us consider a vortex tube of length $\ell$ and cross-rection $A\sim r^{2}$. Under the action of the external mean shear, the vortex tube is stretched and its cross-section decreases. This means enstrophy at a larger scale has been converted to a more intense enstrophy at a smaller scale. Based on this assumption, the transfer of enstrophy among scales in the inertial range is driven by eddies with scale $r$ and velocity $\delta u\sim u_{r}\sim\langle\varepsilon_{f}^{\delta q^{2}}\rangle^{1/3}r^{1/3}$. When dealing with $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$, we are concerned with the scale-by-scale vorticity $\omega_{r}\sim\langle\varepsilon_{f}^{\delta q^{2}}\rangle^{1/3}r^{-2/3}$ to obtain the observed scaling law, i.e.,
@@
$\left\langle{\Phi_{f}^{\delta\omega^{2}}}\right\rangle(r)\sim\frac{1}{S(r)}\int_{S(r)}\text{d}S\ u_{r}\omega^{2}_{r}n_{j}\sim r^{-2}r^{2}\ \left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle r^{1/3}r^{-4/3}=\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle r^{-1}\sim r^{-1}.$ (28)
Refer to captionRefer to captionRefer to caption
@@
Figure 10: (left) Dependence of the Vortex stretching term on $r$ for different $De$. (right) Dependence of $\left\langle{\pi^{\delta\omega^{2}}}\right\rangle/\left\langle{\varepsilon_{f}^{\delta\omega^{2}}}\right\rangle$ on $De$.
We now discuss how the presence of polymers influence the scale-by-scale budget for $\left\langle{\delta\omega^{2}}\right\rangle$. The addition of polymers gives three additional terms in the budget equation, i.e., $\langle\pi^{\delta\omega^{2}}\rangle$ that describes the exchange of enstrophy between the polymeric and fluid phases, and $\langle\Phi_{p}^{\delta\omega^{2}}\rangle$ and $\langle D_{p}^{\delta\omega^{2}}\rangle$ that represent the inertial and dissipative enstrophy flux contribution resulting from the fluid/polymer interaction. In other words, like for $\left\langle{\delta q^{2}}\right\rangle$, an additional source/sink and alternative mechanisms for the enstrophy scale transfer and dissipation arise. For small $De<1$, both $\langle\Phi_{p}^{\delta\omega^{2}}\rangle$ and $\langle\pi^{\delta\omega^{2}}\rangle$ are rather small, and the classic enstrophy production $\langle V_{st}\rangle$ and transfers $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$ are only marginally influenced by the polymers; see the top right and centre left panels of figure 9. For these $De$, $\langle\pi^{\delta\omega^{2}}\rangle<0$, meaning that on average the polymers act as a sink of enstrophy for the fluid phase; see also figure 10. For large Deborah numbers the scenario is different. As $De$ increases, the intensity of $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$ only slightly changes, indicating that the polymers only marginally affect the classical transfer of enstrophy from larger to smaller scales. However, for $De>1$ the intensity of $\langle\Phi_{p}^{\delta\omega^{2}}\rangle$ and $\langle D_{p}^{\delta\omega^{2}}\rangle$ equals that of $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$ and $\langle D_{f}^{\delta\omega^{2}}\rangle$: for large elasticity, the two fluid- and polymer-driven scale-by-scale transfer and dissipative processes have comparable relevance at all scales $r$. This differs from what is seen for $\left\langle{\delta q^{2}}\right\rangle$, where the global energy cascade is dominated by the nonlinear flux at the larger scales and by the polymeric flux at the smaller ones.
The vortex stretching is strongly modulated for $De>1$. Figure 10 shows that $\langle V_{st}\rangle(r)$ largely weakens as $De$ increases, accordingly with the experimental results of Liberzon et al. (2005) and Liberzon et al. (2006). To be quantitative, we measure $\langle V_{st}\rangle/\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle/r|_{r=L/2}\approx 1.1$ for $De=0$ and $\langle V_{st}\rangle/\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle/r|_{r=L/2}\approx 0.35$ for $De=9$. Notably, the weakening of the vortex stretching is accompanied by a change of sign of $\langle\pi^{\delta\omega^{2}}\rangle$, that for $De>1$ becomes positive and balances the viscous dissipation; see figures 4 and 10. In other words, at large $De$ the polymers annihilate the classical production of enstrophy via vortex stretching, but its interaction with the fluid phase results into a new net production of enstrophy at all scales. As discussed in the following section, this change of behaviour at large $De$ can be explained with a change of the local flow topology.
5.2 Local strain
In this subsection, we relate the influence of the polymeric additives to the vortex stretching $V_{st}(r)$ by looking at the local topology of the flow. The local topology of the flow can be described using the three principle invariants of the velocity gradient tensor $A_{ij}\equiv\partial u_{j}/\partial x_{i}$ (Davidson, 2004; Meneveau, 2011). $A_{ij}$ can be decomposed into its symmetric and antisymmetric parts, i.e. the rate-of-strain tensor $S_{ij}$ and the rate-of-rotation tensor $W_{ij}$
$A_{ij}=\frac{\partial u_{i}}{\partial x_{j}}=\underbrace{\frac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right)}_{S_{ij}}+\underbrace{\frac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}-\frac{\partial u_{j}}{\partial x_{i}}\right)}_{W_{ij}}.$ (29)
$A_{ij}$ depends on the three principal rates of strain $\alpha\geq\beta\geq\gamma$ and the three components of the vorticity $\bm{\omega}$ along the principal axes. In other words, the local topology of the flow is entirely determined by (i) the three principal rates of strain $\alpha,\beta$ and $\gamma$, (ii) the magnitude of the vorticity, i.e. the enstrophy $\omega^{2}=\bm{\omega}\cdot\bm{\omega}$, and (iii) the orientation of $\bm{\omega}$ relative to the principal axes of strain. Additionally, the incompressibility condition implies that the sum of the eigenvalues of $S_{ij}$, $\alpha+\beta+\gamma=0$. This also means that at least one eigenvalue is not positive ($\gamma\leqslant 0$), while at least one must be non-negative ($\alpha\geqslant 0$). The statistics of vortex stretching are related to those of the eigenvalues of $S_{ij}$ as (see Davidson, 2004)
$\left\langle{\omega_{i}\omega_{j}S_{ij}}\right\rangle=-4\left\langle{\alpha\beta\gamma}\right\rangle.$ (30)

### v2

@@ 4.1.2 The nonlinear and non Newtonian fluxes
4.2 Scale enstrophy $\left\langle{\delta\omega^{2}}\right\rangle$
4.2.1 Scale-by-scale budget
Figure 10: All contributions in budget equation for $\left\langle{\delta\omega^{2}}\right\rangle$. In order from left to right and from top to bottom, the panels are for $De=0$, $De=1/3$, $De=1$ and $De=9$. For $De=0$, the inset shows the budget with the $y$ axis is logarithmic scale. The black dashed line denotes $r^{-2}$.
This section considers the enstrophy structure function. Unlike energy and helicity, enstrophy is not an inviscid invariant of the Navier–Stokes equations. The budget equation for $\left\langle{\delta\omega^{2}}\right\rangle$, indeed, features a source term $\langle V_{st}\rangle$, that quantifies the amount of enstrophy that is created up to the scale $r$ by means of vortex stretching-like processes. This means that even in Newtonian HIT, the notion of cascade used for energy transfer does not apply here, and most of the enstrophy is generated directly at the small scales. However, as stated by several authors (see for example Davidson et al., 2008), vortex stretching does occur in the inertial range of HIT and hence a net transfer of enstrophy to smaller scales is expected at these $r$.
We plot the terms in the budget equation for $\left\langle{\delta\omega^{2}}\right\rangle$ in figure 10. We start looking at the Newtonian case ($De=0$). At intermediate and large scales, vortex stretching substantially matches the dissipation of enstrophy, i.e., $\langle V_{st}\rangle(r)\approx(4/3)\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle r$, resulting in a $\langle V_{st}\rangle(r)\sim r^{1}$ scaling. In this intermediate range of scales the forcing $\langle F^{\delta\omega^{2}}\rangle$ and dissipation $\langle D_{f}^{\delta\omega^{2}}\rangle$ contributions are negligible, and the flux is subdominant $\langle\Phi_{f}^{\delta\omega^{2}}\rangle\ll\langle V_{st}^{\delta\omega^{2}}\rangle$.
An insightful interpretation of the linear scaling $\langle V_{st}^{\delta\omega^{2}}\rangle\sim r^{1}$ arises from the observation that vortex stretching is predominantly contributed by the smallest scales. This can be seen by the dimensional estimate of the scale-by-scale vorticity $\omega_{r}$
$\omega_{r}\sim\frac{u_{r}}{r}\sim\langle\varepsilon_{f}^{\delta q^{2}}\rangle^{1/3}r^{1/3}r^{-1}=\langle\varepsilon_{f}^{\delta q^{2}}\rangle^{1/3}r^{-2/3},$ (46)
@@ 4.2.1 Scale-by-scale budget
which becomes maximum as $r\to\eta$, so that $\omega_{\eta}\sim\langle\varepsilon_{f}^{\delta q^{2}}\rangle^{1/3}\eta^{-2/3}$; here we have used the $u_{r}\sim\langle\varepsilon_{f}^{\delta q^{2}}\rangle^{1/3}r^{1/3}$ Kolmogorov prediction. Now, to estimate the cumulative vortex stretching $\left\langle{V_{st}}\right\rangle$ up to scale $r$,
@@
$\left\langle{V_{st}}\right\rangle(r)$ $=\frac{1}{\mathcal{S}(r)}\int_{\mathcal{V}(r)}\ \underbrace{2\left(\left\langle{\delta\omega_{i}\omega_{j}^{*}\delta\left(\frac{\partial u_{i}}{\partial x_{j}}\right)}\right\rangle+\left\langle{\delta\omega_{i}\delta\omega_{j}\left(\frac{\partial u_{i}}{\partial x_{j}}\right)^{*}}\right\rangle\right)}_{v_{st,r}}\text{d}\Omega,$ (47)
we begin by estimating the scale-by-scale vortex stretching $v_{st,r}$, as
$v_{st,r}\sim\omega_{r}\omega_{r}\frac{u_{\eta}}{\eta}.$ (48)
@@
We now use the notion that $v_{st,r}$ gets a maximal contribution from the smallest scales (Davidson et al., 2008) and write
$v_{st,\eta}\sim\omega^{2}_{\eta}\ \frac{u_{\eta}}{\eta}\sim\langle\varepsilon_{f}^{\delta q^{2}}\rangle^{2/3}\eta^{-4/3}\langle\varepsilon_{f}^{\delta q^{2}}\rangle^{1/3}\eta^{1/3}\eta^{-1}=\langle\varepsilon_{f}^{\delta q^{2}}\rangle\eta^{-2}\sim\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle,$ (49)
where we have used the $\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle\sim\langle\varepsilon_{f}^{\delta q^{2}}\rangle\eta^{-2}$ dimensional estimate. By neglecting the contribution to vortex stretching from larger scales, we thus estimate the cumulative vortex stretching up to a scale $r$ as $\langle V_{st}\rangle(r)\approx 1/\mathcal{S}(r)\int_{\mathcal{V}(r)}v^{st}_{\eta}\text{d}\Omega\sim r^{-2}r^{3}\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle\sim\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle r$, which is consistent with the data plotted in figure 10.
@@
Moreover, figure 10 shows that the nonlinear flux exhibits a $\langle\Phi_{f}^{\delta\omega^{2}}\rangle(r)\sim r^{-1}$ scaling behaviour at the intermediate scales, with its magnitude increasing as $r$ decreases. Since enstrophy is predominantly generated at the smallest scales, the inter-scale transfer rate intensifies with decreasing $r$, reaching a maximum near the transition between the inertial and dissipative ranges. Following the dimensional analysis approach of Davidson et al. (2008), this scaling can be interpreted by analogy with the vortex stretching mechanism. Consider, for example, a vortex tube of length $\ell$ and cross-sectional area $A\sim r^{2}$. When subjected to external shear, the tube is stretched, reducing its cross-section and thereby increasing its local vorticity. This illustrates how enstrophy at a larger scale is transferred to more intense enstrophy at smaller scales. Based on this picture, the enstrophy transfer in the inertial range is driven by eddies of size $r$, with characteristic velocity $\delta u(r)\sim u_{r}\sim\langle\varepsilon_{f}^{\delta q^{2}}\rangle^{1/3}r^{1/3}$ and vorticity $\delta\omega(r)\sim\omega_{r}\sim\langle\varepsilon_{f}^{\delta q^{2}}\rangle^{1/3}r^{-2/3}$. These scalings naturally lead to the observed behaviour:
@@
$\left\langle{\Phi_{f}^{\delta\omega^{2}}}\right\rangle(r)\sim\frac{1}{\mathcal{S}(r)}\int_{\mathcal{S}(r)}\ u_{r}\omega^{2}_{r}\text{d}\Sigma\sim r^{-2}r^{2}\ \left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle r^{1/3}r^{-4/3}=\left\langle{\varepsilon_{f}^{\delta q^{2}}}\right\rangle r^{-1}.$ (50)
Figure 11: Dependence of $\langle v_{st}\rangle$ and $\langle\pi^{\delta\omega^{2}}\rangle$ on $De$. All quantities are made dimensionless with $\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle$. Here $v_{st}=\omega_{i}\omega_{j}S_{ij}$ is the vortex-stretching production term that appears in the budget equation for $\langle\omega^{2}\rangle$. The bars show the standard deviation computed from six different snapshots.
We now examine the influence of polymers on the budget of $\langle\delta\omega^{2}\rangle$. Their presence introduces two additional terms into the enstrophy budget: a polymeric source/sink term, $\pi^{\delta\omega^{2}}$, and the associated flux term, $\Phi_{p}^{\delta\omega^{2}}$. In other words, analogous to the case of $\langle\delta q^{2}\rangle$, polymers contribute with a source/sink mechanism and an alternative pathway for interscale enstrophy transfer. Recall that for enstrophy we have not split $\Phi_{p}^{\delta\omega^{2}}$ into the inertial and dissipative contributions.
For small Deborah numbers ($De<1$), the polymeric flux $\langle\Phi_{p}^{\delta\omega^{2}}\rangle$ and source term $\langle\pi^{\delta\omega^{2}}\rangle$ remain weak, and the classical enstrophy production $\langle V_{st}\rangle$ and transfer $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$ are only marginally affected by the polymers; see the top right panel of figure 10. In this regime $\langle\pi^{\delta\omega^{2}}\rangle<0$, indicating that polymers act, on average, as a sink of enstrophy for the fluid phase (see also figure 11). Moreover, $\langle\Phi_{p}^{\delta\omega^{2}}\rangle<0$ at all scales, implying that enstrophy is transferred from larger to smaller scales even along the polymer-induced interscale transfer.
At large Deborah numbers, the dynamics of the enstrophy transfer undergoes a marked transformation. As $De$ increases, the average fluid contribution $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$ remains nearly unchanged, indicating that polymers exert limited influence on the classical inertia-driven enstrophy interscale transfer. However, once $De\geq 1$, the polymeric contribution $\langle\Phi_{p}^{\delta\omega^{2}}\rangle$ becomes comparable in magnitude to $\langle\Phi_{f}^{\delta\omega^{2}}\rangle$, and displays a different behaviour across scales. At intermediate scales, both fluid- and polymer-driven transfer processes contribute to the forward enstrophy transfer, from larger to smaller scales, though both remain subdominant compared to the vortex-stretching term. At sufficiently small scales, the scenario changes. For $De\geq 1$, the polymeric flux $\langle\Phi_{p}^{\delta\omega^{2}}\rangle$ becomes negative as $r\rightarrow 0$, suggesting that the fluid–polymer interaction facilitates an inverse enstrophy transfer from smaller to larger scales. Notably, this shift is accompanied by a sign reversal in $\langle\pi^{\delta\omega^{2}}\rangle$ (see §§3 and 5). The data suggest that, at these values of $De$, polymers generate enstrophy at the smallest scales, which is then transferred towards slightly larger scales before being ultimately dissipated by viscosity. This behaviour stands in contrast to that observed for $\langle\delta q^{2}\rangle$, where the nonlinear flux dominates the energy cascade at large scales and the polymeric flux becomes dominant at small scales, but both consistently drive a forward energy transfer across all $r$.
A further key effect at high $De$ is the strong modulation of the vortex-stretching term. Figure 11 illustrates this by showing the dependence of $\langle v_{st}\rangle=\langle\omega_{i}\omega_{j}S_{ij}\rangle$ on $De$, where $S_{ij}=(\partial u_{i}/\partial x_{j}+\partial u_{j}/\partial x_{i})/2$ is the rate-of-strain tensor. The quantity $v_{st}$ represents the one-point vortex-stretching production of enstrophy $\omega^{2}$ (Tsinober, 2001), and satisfies $\langle V_{st}^{\delta\omega^{2}}\rangle\rightarrow 2\langle v_{st}\rangle$ as $r\rightarrow\infty$. As shown in figure 11, $\langle v_{st}\rangle$ decreases significantly with increasing $De$, in agreement with earlier findings by Liberzon et al. (2005, 2006); Cai et al. (2010); ur Rehman et al. (2022). Remarkably, this attenuation of vortex stretching is closely accompanied by the striking sign reversal in $\langle\pi^{\delta\omega^{2}}\rangle$, which becomes negative and approximately balances the viscous dissipation $\langle\pi^{\delta\omega^{2}}\rangle\approx\langle\varepsilon_{f}^{\delta\omega^{2}}\rangle$ for $De\geq 1$. In other words, at high elasticity, polymers not only suppress classical enstrophy production via vortex stretching, but also introduce a different dynamical regime, in which polymer stresses act as a net source of enstrophy.
As discussed in §5, the suppression of vortex stretching and the sign reversal in $\langle\pi^{\delta\omega^{2}}\rangle$ with increasing $De$ are closely linked to changes in the local flow topology and alignments, a transformation that redefines enstrophy dynamics in polymer-laden turbulence.
4.3 Scale helicity
4.3.1 Scale-by-scale budget
Figure 12: All contributions in budget equation for $\left\langle{\delta h}\right\rangle$. In order from top-left to bottom-right, the panels are for $De=0$, $De=1/3$, $De=1$ and $De=9$. The yellow and blue shaded regions identify the inertial and elastic ranges of scales. Here $\Phi_{t}=\Phi_{f}+\Phi_{p}$ and $D_{t}=D_{f}+D_{p}$.
We now move to the helicity structure function $\delta h$. In the present set up, turbulence is sustained with the ABC forcing that injects both energy and helicity at the largest scales. As such, like energy, helicity is expected to cascade from large to small scales in the inertial range of scales, and to eventually vanish at the small scales where the mirror symmetry $\bm{u}\cdot\bm{\omega}=0$ is restored.
Our results, presented in figure 12, provide clear evidence of a dual direct cascade of energy and helicity across all values of $De$, consistent with the theoretical predictions and numerical observations by Kraichnan (1973), Polifke & Shtilman (1989), and Borue & Orszag (1997).
In the purely Newtonian case, helicity is injected at the largest scales ($r>r_{IE}$) at a rate that matches its dissipation, i.e. $\langle F^{\delta h}\rangle=(4/3)\langle\varepsilon_{f}^{\delta h}\rangle r$. It is then transferred from large to small scales within the inertial range ($r_{DI}<r<r_{IE}$), where $\langle\Phi_{f}^{\delta h}\rangle\sim-(4/3)\langle\varepsilon_{f}^{\delta h}\rangle r$, and is ultimately dissipated by viscous effects at the smallest scales.
As with the energy $\langle\delta q^{2}\rangle$, polymers introduce additional helicity transfer and dissipation mechanisms, whose significance increases with $De$. Specifically, the helicity cascade is governed predominantly by the nonlinear flux $\langle\Phi_{f}^{\delta h}\rangle$ in the range $r_{p}^{*}\lessapprox r<r_{IE}$, and by the polymeric flux $\langle\Phi_{p}^{\delta h}\rangle$ in the smaller-scale range $r_{DI}<r\lessapprox r_{p}^{*}$. As $De$ increases and the polymers stretch more, the inertia-driven helicity cascade $\langle\Phi_{f}^{\delta h}\rangle$ is progressively suppressed, mirroring the behaviour observed for $\langle\Phi_{f}^{\delta q^{2}}\rangle$.
The emergence of two distinct cascade mechanisms dominating in separate scale ranges is consistent with the multiscaling behaviour of $\langle\delta h\rangle$ reported in figure 2.
Refer to captionRefer to caption
@@ 4.3.1 Scale-by-scale budget
Figure 13: Dependence of $\langle\Phi_{f,a}^{\delta h}\rangle$ (left) and $\langle\Phi_{f,b}^{\delta h}\rangle$ (right) on $r$ for different DeDe.
Figure 13 offers a more detailed view of how polymer additives influence the nonlinear helicity flux $\langle\Phi_{f}^{\delta h}\rangle$. As discussed in §2, this flux comprises two distinct contributions. The first is the classical convective term, $\langle\Phi_{f,a}^{\delta h}\rangle$, associated with the correlation $\langle\delta h\delta u_{j}\rangle$. The second, $\langle\Phi_{f,b}^{\delta h}\rangle$, captures a transfer mechanism involving the correlation $\langle\delta\omega_{j}\delta q^{2}\rangle$. In the Newtonian case, figure 13 shows that $\langle\Phi_{f,a}^{\delta h}\rangle$ clearly dominates, being nearly twice as large as $\langle\Phi_{f,b}^{\delta h}\rangle$ across all scales. This indicates that, in Newtonian HIT, helicity is primarily transferred through the convective mechanism. As $De$ increases, however, both components of the flux are progressively suppressed in the elastic range of scales, reflecting the polymers’ ability to attenuate both transfer pathways while redirecting a portion of the helicity flux through the polymeric channel. Notably, polymers have a stronger damping effect on $\langle\Phi_{f,a}^{\delta h}\rangle$ than on $\langle\Phi_{f,b}^{\delta h}\rangle$. As a result, at $De=9$, the two contributions become comparable on average, with $\langle\Phi_{f,a}^{\delta h}\rangle\approx\langle\Phi_{f,b}^{\delta h}\rangle$.
4.3.2 Relative helicity
Refer to captionRefer to caption
Figure 14: (left) Dependence of the relative helicity $r\left\langle{\delta h}\right\rangle/(2\left\langle{\delta q^{2}}\right\rangle)$ on $r$ for different DeDe. (right) Dependence of $\left\langle{\delta h}\right\rangle/\left\langle{\delta q^{2}}\right\rangle$ on $r$ for different DeDe.
We now examine the relative helicity, defined as $h_{r}\equiv r\langle\delta h\rangle/(2\langle\delta q^{2}\rangle)$ following Borue & Orszag (1997), to assess how polymers influence the flow’s tendency to restore mirror symmetry at small scales. As shown in figure 14, the flow exhibits maximal helicity at large scales, with $h_{r}$ decaying monotonically as $r$ decreases, indicating that helicity becomes dynamically negligible at small scales.
In the Newtonian case, we recover the classical scaling $h_{r}\sim r^{1}$ within the inertial range, consistent with the findings of Borue & Orszag (1997). This behaviour is illustrated in the right panel of figure 14, where the ratio $\langle\delta h\rangle/\langle\delta q^{2}\rangle\approx 0.75$ is observed, comparable to that found for passive scalar transport in isotropic turbulence (Borue & Orszag, 1997).
In PHIT, however, the decay of $h_{r}$ with $r$ is noticeably slower in the intermediate range, and the Newtonian $r^{1}$ scaling in the inertial range is lost. This reflects the fact that $\langle\delta h\rangle$ and $\langle\delta q^{2}\rangle$ scale differently for $De>0$, as previously shown in figure 2. Furthermore, we find that $h_{r}$ increases systematically with $De$ across all scales, suggesting a growing asymmetry with increasing elasticity.
At the smallest scales, the ratio $\langle\delta h\rangle/\langle\delta q^{2}\rangle$ tends to a plateau for all cases, which reflects the smoothness of the velocity and vorticity fields and is consistent with a Taylor expansion representation (see also figure 2). Overall, the results in figure 14 indicate that polymers act to break mirror symmetry, and that this symmetry-breaking effect becomes more pronounced with increasing elasticity.
5 Local structure of the flow
In this section, we examine how polymer additives modify the local flow topology. While the preceding section has focused on the budgets of energy, enstrophy, and helicity, the local structure of the flow provides a complementary perspective, that connects small-scale dynamics more directly to the mechanisms of production, transfer and dissipation of the three quantities. Specifically, we relate the polymer-induced changes in energy transfer and enstrophy production to alterations in local strain and rotation dynamics.
The flow topology is characterised using the three principal invariants of the velocity gradient tensor $A_{ij}=\partial u_{j}/\partial x_{i}$ (Davidson, 2004; Meneveau, 2011), which offer compact yet insightful descriptors of the local kinematics.
5.1 Local straining state and alignments
The velocity gradient tensor $A_{ij}=\partial u_{j}/\partial x_{i}$ can be decomposed into its symmetric and antisymmetric components, namely $S_{ij}$ and the rate-of-rotation tensor $W_{ij}$:
$A_{ij}=\frac{\partial u_{j}}{\partial x_{i}}=\underbrace{\frac{1}{2}\left(\frac{\partial u_{j}}{\partial x_{i}}+\frac{\partial u_{i}}{\partial x_{j}}\right)}_{S_{ij}}+\underbrace{\frac{1}{2}\left(\frac{\partial u_{j}}{\partial x_{i}}-\frac{\partial u_{i}}{\partial x_{j}}\right)}_{W_{ij}}.$ (51)
The tensor $S_{ij}$ admits a set of real eigenvalues $\alpha\geq\beta\geq\gamma$, corresponding to the principal strain rates, with associated orthonormal eigenvectors. The local kinematics of the flow can thus be characterised by the triplet $(\alpha,\beta,\gamma)$, by the orientation of the vorticity vector $\bm{\omega}$ with respect to the eigenvectors of $S_{ij}$, and by the magnitude of the vorticity vector, i.e. the enstrophy. In incompressible flows, the trace of $S_{ij}$ vanishes, yielding the constraint $\alpha+\beta+\gamma=0$. This implies that at least one eigenvalue must be non-positive ($\gamma\leq 0$), and at least one must be non-negative ($\alpha\geq 0$).
The vortex stretching term is related to the eigenvalues of $S_{ij}$ through the identity (see Davidson, 2004):
$\langle v_{st}\rangle=\langle\omega_{i}\omega_{j}S_{ij}\rangle=-4\langle\alpha\beta\gamma\rangle.$ (52)